# Surprisingly High Redundancy in Electronic Structure Data

### v1

@@ II.1 Aluminum
Figure 4: Aluminum system: (a) The error in the electron density prediction for various pruning factors. The shaded region represents the range (maximum to minimum) of three ML models around the mean, shown by a solid line. (b) The error in energy obtained from the predicted electron density for various pruning factors.
Figure 5: Aluminum system at 1500K. Electron density obtained by: (a) DFT and (b-d) ML model for original, 90% CCS and 99% CCS. (e-h) A two-dimensional (2D) slices of the simulation cell corresponding to Figures (a-d) showing electron density. (i) Electron density along the solid line, shown in Figures (e-h). (j-l) Absolute error in ML-predicted electron density for original, 90% CCS, and 99% CCS corresponding to Figures (f-h). The unit of electron density is $\text{e}\cdot\text{Bohr}^{-3}$, where e denotes the electronic charge.
The errors (NRMSE) in electron density predictions for the three pruning approaches — random, GraNd, and CCS — are plotted against various pruning factors in Figure 4(a). GraNd, random, and CCS perform nearly as well as the original dataset up to $40\%$, $80\%$, and $90\%$ pruning, respectively. The performance of the GraNd technique deteriorates more rapidly than random and CCS techniques. The exceptional performance of CCS at higher pruning rates, such as $80\%$ and $90\%$, is evident, with the increase in NMRSE compared to the original dataset being only $0.7\%$ and $2.1\%$, respectively. CCS also performs surprisingly well at $99\%$ pruning (i.e. two orders of magnitude less data), showing only a nominal increase in NMRSE.
To assess the quality of the prediction further, the total energy of the system is calculated from the ML-predicted electron density, via postprocessing. The total energy error (Hartree/atom) relative to the KS-DFT calculations, are shown in Figure 4(b). This is essential in understanding if it is possible to maintain chemical accuracy in the energies computed from the electron density predictions of ML models trained on pruned data. All three methods maintain near-original accuracy in the energy up to $80\%$ pruning. In particular, all three techniques maintain chemical accuracy even at $99\%$ pruning, with the CCS and random pruning techniques surpassing the accuracy of the GraNd technique at such high pruning fractions.
To compare the accuracy of the ML models against KS-DFT over the entire spatial domain, we plot the ML predicted electron density fields as obtained using original dataset, the $90\%$ CCS dataset, and the $99\%$ CCS dataset, in Figure 5(a-h). All three ML models show excellent agreement with the DFT calculations. The error fields of the ML model predictions with respect to the KS-DFT simulations are shown in Figure 5(j-l), which confirms that the error for $90\%$ or $99\%$ CCS does not show significant increase compared to the original ML model. The electron density along the solid line shown in figures (e-h) is plotted in Figure  5(i). The details of the 3D error fields is also shown via iso-surfaces in the supplemental materials V.8, Figure 21, which shows negligible increase in error for $90\%$ and $99\%$ pruned data.
II.2 CrFeCoNi Alloy
We find that just $10\%$ of the data, selected through the CCS method, can match the performance of the original dataset while $1\%$ of the data can be selected to yield a exceptionally accurate ML model, that maintains chemical accuracy in the energies.
@@ II.2 CrFeCoNi Alloy
Figure 6: CrFeCoNi system: (a) Error in the electron density prediction for various pruning factors. (b) Error in energy obtained from the predicted electron density for various pruning factors.
Figure 7: $\text{Cr}_{\text{25}}\text{Fe}_{\text{25}}\text{Co}_{\text{25}}\text{Ni}_{\text{25}}$ alloy system at $5000$ K: Electron density obtained by: (a) DFT and (b-d) ML model trained on the original dataset, 90% CCS and 99% CCS based pruned dataset. (e-h) Two-dimensional (2D) slices of the simulation cell corresponding to Figures (a-d). (i) Electron density along the solid line, shown in Figures (e-h). (j-l) Absolute error in ML-predicted electron density for original dataset, 90% CCS-based pruned dataset, and 99% CCS-based pruned dataset. The unit of electron density is $\text{e}\cdot\text{Bohr}^{-3}$, where e denotes the electronic charge.
The errors (NRMSE) in electron density predictions for three pruning approaches — random, GraNd and CCS — are plotted against various pruning factors in Figure 6(a). Random, GraNd, and CCS methods perform nearly as accurately as the original dataset up to $40\%$, $60\%$ and $90\%$ pruning, respectively. The performance of GraNd-based pruning drops catastrophically at higher pruning rates, far more rapidly than random pruning or CCS. Indeed, the exceptional performance of the CCS method is evident even at higher pruning rates, such as $80\%$ and $90\%$, with the variations in errors being less than $1\%$ as compared to the original dataset. CCS also performs very well at $99\%$ pruning (i.e. two orders of magnitude less data), showing a slight increase in the error (NRMSE).
The quality of ML-predicted electron densities are further assessed through post-processing calculations to obtain the total energy of the system. The total energy error (Hartree/atom) relative to KS-DFT calculations are shown in Figure 6(b). This is essential in understanding if it is possible to maintain chemical accuracy in energies from the electron density predictions based on the pruned datasets. All methods maintain near-original accuracy in the energy up to $60\%$ pruning. In particular, the CCS approach provides near original accuracy up to $99\%$ while maintaining chemical accuracy in the energies, and surpassing both random pruning and GraNd, even at such a high pruning rate. Notably, the fluctuations in the energy errors of the original, $90\%$ and $99\%$ pruned datasets are consistent with observations made in ML predictions of the electronic structure of quaternary systems in our recent work [28].
An unscaled version of Figure 6 can be found in the supplemental materials V.3, Figure 14. To compare the accuracy of the ML models against the KS-DFT over the spatial domain, we plot the ML-predicted electron density fields obtained using the original dataset, $90\%$ CCS dataset and $99\%$ CCS dataset, in Figure 7. All three models show remarkable agreement with the DFT calculations. The absolute error in the fields of ML predictions with respect to the KS-DFT calculations are plotted in Figure 7(j-l). The errors in $90\%$ CCS pruned dataset (see Figure 7(k)) are very much similar to the errors in the original dataset(see 7(j)). With $99\%$ pruning, (see Figure 7(l)), the error intensifies slightly throughout the slice. The electron density along the solid line shown in Figure 7(e-h) is plotted in the Figure 7(i). The details of the 3D error fields is also shown via iso-surfaces in the supplemental materials V.8, Figure 23, which shows negligible increase in error for $90\%$ and slight increase for $99\%$ pruned data.
II.3 Water
Figure 9: Water system at 600K: Electron density obtained by: (a) DFT and (b-d) ML model trained on the original dataset, 90% CCS and 99% CCS based pruned dataset. (e-h) Two-dimensional(2D) slices of the simulation cell corresponding to figures (a-d). (i) Electron density along the solid line, shown in Figures (e-h). (j-l) Absolute error in ML-predicted electron density for the original dataset, 90% CCS-based pruned dataset, and 99% CCS-based pruned dataset. The unit of electron density is $\text{e}\cdot\text{Bohr}^{-3}$, where e denotes the electronic charge.
We find that just 10% of the data for the water systems, selected through CCS, can match the performance of the original dataset while 1% of the data can be selected using CCS approach to obtain a very accurate ML model, that maintains chemical accuracy.
The errors (NRMSE) in the three pruning approaches- random, GraNd and CCS are plotted in the Figure 8(a). GraNd, random and CCS performs nearly as accurate as the original dataset up to 60%, 60% and 90% pruning. The performance of the GraNd pruning drops rapidly as compared to the random and CCS methods, which is consistent with the other systems studied here. The exceptional performance of CCS at higher pruning rates, such as 80% and 90% is evident with the increase in error compared to the original dataset being only 23% and 30% respectively. CCS also performs surprisingly well for 99% pruning, but showing a modest increase in error.
The quality of the predictions are further assessed through the post-processing calculations of the energy of the system, obtained from the ML-predicted electron density. The total energy error(Ha/atom) relative to KS-DFT calculations is shown in Figure 8(b). This allows us to quantify the quality of the ML-predicted electron density for all the pruning factors and is essential for understanding if it is possible to maintain chemical accuracy in energies from ML-predicted electron densities that were trained on pruned datasets. All methods maintain near-original accuracy in the energy up to $40$% pruning maintaining chemical accuracy levels. In particular, the random pruning and CCS approach provide near original accuracy up to $90$% while maintaining chemical accuracy level in the energies surpassing GraNd at such high pruning rate. An unscaled version of Figrue 8 is provided in the Supplemental Materials V.3, Figure 15.
To compare the accuracy of the ML models against the KS-DFT over the spatial domain, we plot the ML-predicted electron density fields obtained using the original dataset, 90% CCS dataset, and 99% CCS dataset in Figure 9. All three models show remarkable agreement with the DFT calculations. The absolute error in the fields of ML predictions with respect to the KS-DFT calculations are plotted in Figure 9(j-l). The errors in the 90% CCS pruned dataset (see Figure 9(k)) are very similar to the errors in the original dataset (see Figure 9(j)). With 99% pruning (see Figure 9(l)), the error intensifies slightly throughout the slice. The electron density along the solid line shown in Figure 9(e-h) is plotted in Figure 9(i). The details of the 3D error fields are also shown via iso-surfaces in supplemental materials V.8, Figure 25, which shows a negligible increase in error for 90% pruning and a slight increase for 99% pruned data.
II.4 Generalizability of the ML models
@@
Figure 10: Model generalizability to unseen defect configurations: (Top) Mono-vacancy defect for Aluminum 108 atoms systems. Two-dimensional(2D) slices of electron density obtained by: (a) DFT and (b-d) ML model trained on the original dataset, 90% CCS and 99% CCS based pruned dataset. (e) Electron density along the solid line, shown in Figures (a-d). (Bottom) Mono-vacancy defect for CrFeCoNi 32 atoms systems. 2D slices of electron density obtained by: (f) DFT and (g-i) ML model trained on the original dataset, 90% CCS and 99% CCS based pruned datasets. (j) Electron density along the solid line, shown in Figures (f-i). The unit of electron density is $\text{e}\cdot\text{Bohr}^{-3}$, where e denotes the electronic charge.
@@
The limited generalization capability of ML models beyond their training data remains a fundamental issue of the field. In particular, the generalization capability of ML models developed using pruned data is yet unexplored. To address this issue, we evaluate the generalization capabilities of the ML models developed here, by assessing their performance on a diverse set of systems that differ significantly from, and were not included in, the training data. These test cases include systems with structural defects such as mono- and di-vacancies, even though the models were trained exclusively on defect-free systems.
The ML-predicted densities via CCS ($90$% and $99$% pruning) exhibit remarkable agreement with KS-DFT across all test systems, including both aluminum and CrFeCoNi alloys. Moreover, the prediction errors for defective systems, e.g. mono-vacancy, are comparable in magnitude to those observed for pristine systems, as illustrated in Figure 10. Importantly, the generalization capability of the ML models remains unaffected by upto two orders of magnitude CCS-based pruning. Additional comparisons between ML-predicted and KS-DFT-computed electron density fields for various pruning factors, including the di-vacancy defect for the aluminum system and the CrFeCoNi system, are presented in the Supplemental Material V.5.
@@ III Discussion
We have explored the hidden redundancy in electronic structure datasets for diverse material systems. To the best of our knowledge, this marks the first exploration of redundancy in electronic structure datasets through systematic pruning approaches. In addition to random pruning, we have employed two ML-based pruning techniques, namely GraNd [65] and CCS [68]. GraNd based pruning selects a subset of data having the highest importance scores. The CCS approach selects a subset that ensures comprehensive coverage of the dataset based on a probability density based data distribution coverage. These two ML based pruning methods are compared against random grid-point wise pruning — with the goal of revealing the aforementioned redundancy, and helping select the minimal essential dataset without compromising accuracy. Overall, we are able to identify a large fraction of the electronic structure data that is redundant and which can be eliminated from ML training, without compromising the accuracy or the smoothness of the ML predicted fields.
Random grid-point-wise pruning turns out to be an effective strategy, aligning with findings in the computer vision literature. Although its error rate increases at higher pruning levels, it can outperform several ML-based pruning methods [77, 68].
For all three material systems, at lower pruning rates, GraNd captures the redundancy of the dataset effectively, closely matching or outperforming the original dataset. However, at higher pruning rates, GraNd’s performance drops catastrophically. This occurs because, at large pruning factors, most of the easy-to-learn data points are removed, significantly altering the data distribution and leading to poor data coverage.
@@
We also demonstrate that $90$% and $99$% CCS based pruned datasets retain significant generalization capability of the model, through tests on mono- and divacancies.
In addition to identifying the essential subset of the data, pruning can also accelerate ML model training due to the reduced size of the subset. We found that for the materials considered here, ML model training time on the pruned datasets can be lower by $50\%-80\%$, as compared to the training time on unpruned data (See Section V.7 of the supplemental material). Thus, such pruned datasets can be used to accelerate hyperparameter tuning. In turn, these advancements can be critical in developing ML models that work across chemical space.
@@ IV Methods
The electron density data needed for training, validation, and testing of the ML models was produced via Kohn-Sham Density Functional Theory (KS-DFT) calculations [5, 78] as implemented in the finite-difference based SPARC code [79, 80, 81]. Norm conserving (ONCV) pseudopotentials [82, 83, 84] and GGA PBE exchange-correlation [85] were employed. Prior to using the pseudopotentials for electronic structure data generation, tests were performed to verify that lattice parameters and other physical quantities predicted were consistent with the literature and that the energies and forces were sufficiently converged with respect to the mesh fineness. The mesh spacing chosen for the aluminum, CrFeCoNi, and water systems were $0.25$, $0.20$, and $0.25$ Bohr respectively. A tolerance of $10^{-6}$ was chosen for the self-consistent field (SCF) iteration convergence, and the Periodic-Pulay scheme [86] was employed for convergence acceleration. The KS-DFT calculations for ML data generation only sampled the gamma point in reciprocal space, as is common in the simulation of large condensed matter systems.
@@
While the aforementioned calculations provided the electron densities for a given atomic configuration, additional techniques were employed to efficiently generate a large sample set of atomic configurations. For the aluminum system, Ab Initio Molecular Dynamics (AIMD) simulations were employed, and the electron density data needed for the ML models was extracted from AIMD “snapshots” at regular simulation intervals [27, 19]. A standard Nosé Hoover thermostat [87] with temperatures between $298$ K and $2100$ K was employed, and with Fermi-Dirac smearing at a fixed electronic temperature of $631.554$ K was used. As pointed out in our recent contributions, use of thermalized systems to sample the configuration space leads to more generalizable ML models of the electronic structure [27, 28], due to better sampling of the input features space.
@@
While AIMD simulations are convenient for data generation, they are computationally costly. In order to reduce the data generation expense for the water and CrFeCoNi alloy systems, atomic configurations were generated through classical Molecular Dynamics (MD) simulations employing machine learning based interatomic potentials [28]. No accuracy is sacrificed by this method choice since the electron densities corresponding to each configuration are still produced from high-fidelity KS-DFT calculations. At the same time, by running these classical MD simulations at elevated temperatures and over long times, a wide variety of the atomic configurations can be conveniently generated by sampling the resulting MD trajectories. We employed the Materials Graph Library (MatGL) python package [88, 89], which includes the Materials 3-body Graph Network (M3GNet) — a universal machine-learned interatomic potential — for these purposes. The NVT ensemble with Nosé Hoover thermostat was employed for these MD runs as well. The CrFeCoNi alloy data was collected from high-temperature simulations between $2500$ K and $7000$ K, while the water data was obtained from simulations with temperatures between $300$ K and $600$ K.

### v2

@@ II.2 Comparative assessment of pruning strategies
Figure 5: The error in the electron density prediction for various pruning factors for (a) Aluminum (b) Water (c) SiGeSn (d) CrFeCoNi. The shaded region represents the range (maximum to minimum) of three ML models around the mean, shown by a solid line.
@@
Figure 6: The error in the total energy obtained via postprocessing from the predicted electron density for various pruning factors for (a) Aluminum (b) Water (c) SiGeSn (d) CrFeCoNi.
In the following, we study the performance of three prominent pruning techniques — random, GraNd and CCS — for various pruning factors. ML models for each pruning factor and pruning strategy were trained three times, and the results are reported accordingly. The errors (NRMSE) in electron density predictions for these three pruning approaches are plotted against various pruning factors across all four materials systems in Figure 5. The ML models exhibit minor random fluctuations in their predictions, which is expected due to inherent randomness in the dataset and the model. All three methods perform well at smaller pruning factors across materials, maintaining the accuracy of the original dataset. However, the GraNd method fails catastrophically, with errors rising sharply beyond 40–60% pruning across different materials. Random pruning shows a steady increase in error with pruning factor, preserving the accuracy of the original dataset up to 40–80% and exhibiting rapid error growth beyond that. In contrast, the CCS method outperforms both GraNd and random pruning, maintaining the accuracy of the original dataset up to 80–90% pruning and showing a slower increase in error than random beyond 90%. Remarkably, CCS continues to deliver accurate results at 99% pruning across all materials, with only a modest increase in error compared to the original dataset. The corresponding errors in the total energy of the system, as obtained from post-processing the ML-predicted electron density, are plotted against various pruning factors in Figure 6. For each pruning factor and and strategy, the predicted electron density from one of the three trained ML models was post-processed. All three methods predict the energy nearly as accurately as the original dataset for smaller pruning factors, up to 40%. However, for aluminum all three methods are chemically accurate until 99% pruning, despite a sharp jump in error for the GraNd method beyond 80%. The GraNd method retains chemical accuracy up to 40% pruning for all materials; however, at higher pruning factors, its energy error increases very rapidly for all materials. Random pruning retains chemical accuracy up to 60% pruning for all materials. It maintains chemical accuracy for simple materials, such as aluminum and water, up to 99%, whereas for complex alloys, the limit decreases to 90% for ternary systems and 60% for quaternary systems. The CCS method consistently outperforms the other two approaches, nearly preserving the accuracy of the original dataset up to 90% pruning across all materials. Remarkably, it continues to yield chemically accurate energy predictions even at 99% pruning. Notably, unlike the pure materials considered here — bulk water and aluminum metal — the errors in energies computed for complex alloy systems appear to exhibit random fluctuations with respect to pruning factor for high pruning factors. Thus, for such cases, the errors in the energy, although related with the electron density errors, are not necessarily monotonic with respect to pruning factors. This is consistent with our recent observations for ML based electronic structure prediction of complex quaternary alloys, whereupon the NRMSE in electron density errors alone was found to be not sufficient in completely determining the energy errors [pathrudkar2024HEA]. Nevertheless, we observe that for CCS based pruning, at very high pruning factors, the quaternary systems do achieve chemical accuracy in the energies, while all the other systems achieve significantly higher accuracies. Further details on these pruning results are provided in the supplemental materials (section V.7.)
II.3 Generalizability of the ML models
@@
Figure 7: Model generalizability to unseen configurations: 2D slices of electron density obtained by KS-DFT, ML model trained on the original dataset, 90% CCS, and 99% CCS based pruned datasets. The electron density($\rho$) along the solid line is also compared with KS-DFT, showing remarkable agreement. The energy errors(Ha/atom) are also shown for each case. (Top) Mono-vacancy defect for 32-atoms aluminum systems. (Middle) 216-atoms SiGeSn checkerboard systems (Bottom) Di-vacancy defect for 32-atoms CrFeCoNi systems. The unit of electron density is $\text{Bohr}^{-3}$ in atomic units.
@@
The limited generalization capability of ML models beyond their training data remains a fundamental issue of the field. In particular, the generalization capability of ML models developed using pruned data is yet unexplored. To address this issue, we evaluate the generalization capabilities of the ML models developed here, by assessing their performance on a diverse set of systems that differ significantly from, and were not included in, the training data. These test cases include systems with structural defects such as mono-vacancies, di-vacancies, and handcrafted checkerboard alloy systems featuring species segregation. We find that the generalization capability of the ML models remains unaffected for up to 99% CCS-based pruning.
The ML-predicted densities via CCS (9090% and 9999% pruning) exhibit remarkable agreement with KS-DFT across all of the above test systems, as illustrated in Figure 7. We further compute each system’s energy from ML-predicted electron densities obtained for the original dataset, 90% CCS and 99% CCS based pruned datasets, as illustrated in Figure 7 (see Supplemental Materials, section V.5 for further details). For $99\%$ pruning, we observe chemically accurate predictions for all cases except for the di-vacancies in the quaternary system, where the error is still relatively low ($\mathcal{O}(10^{-3})$ Ha/atom). Additional comparisons between ML-predicted and KS-DFT-computed electron density for various pruning factors, including mono-vacancy defects for the CrFeCoNi system, are presented in the supplemental material(section V.5).
@@ III Discussion
We have explored the hidden redundancy in electronic structure datasets for diverse material systems. To the best of our knowledge, this marks the first exploration of redundancy in electronic structure datasets through systematic pruning approaches. In addition to random pruning, we have employed two ML-based pruning techniques, namely GraNd [paul2021deep] and CCS [zheng2023_CCS]. GraNd based pruning selects a subset of data having the highest importance scores. The CCS approach selects a subset that ensures comprehensive coverage of the dataset based on a probability density-based coverage strategy. These two ML based pruning methods are compared against random grid-point wise pruning — with the goal of revealing the aforementioned redundancy, and helping select the minimal essential dataset without compromising accuracy. Overall, we are able to identify a large fraction of the electronic structure data that is redundant and which can be eliminated from ML training, without compromising the accuracy or the smoothness of the ML predicted fields.
Random grid-point-wise pruning turns out to be an effective strategy, aligning with findings in the computer vision literature. Although its error rate increases steadily, it can outperform several ML-based pruning methods at higher pruning factors  [guo2022deepcore, zheng2023_CCS].
For all four material systems, at lower pruning rates, GraNd captures the redundancy of the dataset effectively, closely matching or outperforming the original dataset. However, at higher pruning rates, GraNd’s performance drops catastrophically. This occurs because, at large pruning factors, most of the easy-to-learn data points are removed, significantly altering the data distribution and leading to poor data coverage.
@@
We also demonstrate that 9090% and 9999% CCS based pruned datasets retain significant generalization capability of the model, through tests on mono-vacancies, di-vacancies and handcrafted systems that were not included in training. We also compute the system’s energy from the ML-predicted electron densities and demonstrate that the 90% and 99% CCS based pruned datasets can provide chemically accurate predictions.
In addition to identifying the essential subset of the data, pruning can also accelerate ML model training due to the reduced size of the subset. We found that for the materials considered here, ML model training time on the $99\%$ pruned datasets can be lower by $75\%-90\%$, as compared to the training time on the original data (See Section V.6 of the supplemental material). Thus, such pruned datasets can be used to accelerate hyperparameter tuning. In turn, these advancements can be critical in developing ML models that work across chemical space.
@@ IV Methods
The electron density data needed for training, validation, and testing of the ML models was produced via Kohn-Sham Density Functional Theory (KS-DFT) calculations [kohn1965self, Martin_ES] as implemented in the finite-difference based SPARC code [xu2021sparc, xu2020m, ghosh2017sparc]. Norm conserving (ONCV) pseudopotentials [hamann2013optimized, schlipf2015optimization, van2018pseudodojo] and GGA PBE exchange-correlation [perdew1996generalized] were employed. Prior to using the pseudopotentials for electronic structure data generation, tests were performed to verify that lattice parameters and other physical quantities predicted were consistent with the literature and that the energies and forces were sufficiently converged with respect to the mesh fineness. The mesh spacing chosen for the aluminum, water, SiGeSn and CrFeCoNi systems were $0.25$, $0.25$, $0.40$ and $0.20$ Bohr respectively. A tolerance of $10^{-6}$ was chosen for the self-consistent field (SCF) iteration convergence, and the Periodic-Pulay scheme [banerjee2016periodic] was employed for convergence acceleration. The KS-DFT calculations for ML data generation only sampled the gamma point in reciprocal space, as is common in the simulation of large condensed matter systems.
@@
While the aforementioned calculations provided the electron densities for a given atomic configuration, additional techniques were employed to efficiently generate a large sample set of atomic configurations. For the aluminum and SiGeSn alloy systems, Ab Initio Molecular Dynamics (AIMD) simulations were employed, and the electron density data needed for the ML models was extracted from AIMD “snapshots” at regular simulation intervals [pathrudkar2024electronic, zepeda2021deep]. A standard Nosé Hoover thermostat [evans1985nose] with temperatures between $298$ K and $2100$ K was employed for the aluminum systems, while a temperature of $2400$ K was employed for the SiGeSn systems. Fermi-Dirac electronic smearing at a fixed electronic temperature of $631.554$ K was used. As pointed out in our recent contributions, use of thermalized systems to sample the configuration space leads to more generalizable ML models of the electronic structure [pathrudkar2024electronic, pathrudkar2024HEA], due to better sampling of the input features space.
@@
Although AIMD simulations are convenient for data generation, they are computationally costly. In order to reduce the data generation expense for the water and CrFeCoNi alloy systems, atomic configurations were generated through classical Molecular Dynamics (MD) simulations employing machine learning based interatomic potentials [pathrudkar2024HEA]. No accuracy is sacrificed by this method choice since the electron densities corresponding to each configuration are still produced from high-fidelity KS-DFT calculations. At the same time, by running these classical MD simulations at elevated temperatures and over long times, a wide variety of the atomic configurations can be conveniently generated by sampling the resulting MD trajectories. We employed the Materials Graph Library (MatGL) python package [chen2021learning, chen2022universal], which includes the Materials 3-body Graph Network (M3GNet) — a universal machine-learned interatomic potential — for these purposes. The NVT ensemble with Nosé Hoover thermostat was employed for these MD runs as well. The CrFeCoNi alloy data was collected from high-temperature simulations between $2500$ K and $7000$ K, while the water data was obtained from simulations with temperatures between $300$ K and $600$ K.